# ESIGMAHM: An Eccentric, Spinning inspiral-merger-ringdown waveform model with Higher Modes for the detection and characterization of binary black holes

### v1

@@ V Importance of including eccentricity and higher-order waveform modes
An important effect that immediately distinguishes an eccentric binary from a quasi-circular one is the increased rate of inspiral, leading to a shorter time to merger. We can quantify this by computing the number of GW cycles of the dominant-mode from a given frequency to merger, for both eccentric and equivalent quasi-circular cases (same black hole masses and spins). We compute the differences between them
@@
starting from a GW frequency of $10$Hz. Positive values of $\Delta\rm{N}_{\rm{GW}}$ imply a shorter eccentric signal. We show $\Delta\rm{N}_{\rm{GW}}$ in Fig. 9 as a function of total mass and mass ratio $M-q$ (left panel), and eccentricity and mass ratio $e-q$ (middle and right panels). The component spins are fixed to moderate values of $(0.5,0.5)$ in all panels, and orbital eccentricity is held constant at $0.05$ in the left panel. The middle panels correspond to a fixed binary mass of $20M_{\odot}$, while the right panel is for heavier binaries with mass $60M_{\odot}$. In the left panel showing $M-q$ plane, we see that there is a sharp increase in $\Delta\rm{N}_{\rm{GW}}$ with increasing mass ratio for small total masses, reaching as high as $(\sim 37.1)$ fewer GW cycles for $10M_{\odot}$ eccentric binaries with mass ratios $q\simeq 8$ even for relatively small value of eccentricity at $10$Hz $(e_{10})$ of $0.05$. For masses larger than $20M_{\odot}$ the difference is relatively smaller $(\lesssim 4)$ irrespective of mass ratio. The mass dependence is also relatively weak indicating that binaries starting with $e_{10}=0.05$ with masses above $20M_{\odot}$ rapidly circularize in the band of ground-based detectors. In $e-q$ plane shown in the middle panel, we see clearly a stronger non-linear growth in $\Delta\rm{N}_{\rm{GW}}$ as we increase initial eccentricity or (to a lesser extent) the binary mass ratio. For these $20M_{\odot}$ binaries, $\Delta\rm{N}_{\rm{GW}}$ grows rapidly and reaches $\mathcal{O}(10^{2})$ cycles when eccentricities is larger than $0.25$ for $q=2$ binaries, and when $e_{10}>0.15$ for $q=6$ systems. Compare this panel with the right panel to note the effect of increasing binary mass from $20$ to $60M_{\odot}$, a range which spans the masses of a bulk of sources observed by the current generation detectors so far [6, 7, 9, 185]. At $60M_{\odot}$ we notice a smaller yet substantial change in the number of GW cycles from eccentric binaries compared to quasi-circular ones. This is consistent with there being much fewer signals cycles in frequency band for higher masses. Note that the left and middle panels explore different parts of the parameter space, with only a horizontal strip of constant $M=20M_{\odot}$ in the left panel being common with the $e_{10}=0.05$ horizontal strip in the middle panel.
@@
Figure 12: Effect of eccentricity and higher-order modes are shown here on matched-filtering based GW searches. We show the distribution of fitting factors for different kinds of signal populations (as labeled in the legend), against a template bank of quasi-circular aligned-spin BBH waveform templates, similar to what is used in contemporary GW searches in LIGO-Virgo-KAGRA data. Comparing the blue and green curves shows the effectiveness of the same template bank in recovering eccentric vs quasi-circular BBH merger GW signals. This shows how currently used template banks will lose a sizable fraction of signal information from eccentric sources. The effect gets worse when sub-dominant harmonics are included in signal waveforms (comparing light and dark blue curves).
@@
Figure 13: Effect of eccentricity and higher-order waveform modes are shown on the detectability of eccentric binary mergers by current LIGO-Virgo-KAGRA searches. We show fitting factors for eccentric dominant-mode signals in the top row, and for eccentric multi-modal signals in the bottom row. In each panel, the same is shown as a function of source eccentricity at $10$Hz, and either binary mass ratio (left column) or total mass (right column). Component masses are sampled uniformly between $10-50M_{\odot}$ and spins uniformly between $[-0.9,0.9]$. Even though fitting factor values go down to 0.7 or so (as seen in Fig. 12), we set the lower limit of these panels’ color bars to 0.9 in order to improve the visual contrast and highlight those parameter space regions where fitting factors remained above 0.97. We immediately note that for realistic signals, i.e. the bottom row, our current GW searches that employ only dominant-mode aligned-spin quasi-circular waveform templates, we will lose more than $10\%$ of signal information (optimal SNR) for sources with mass-ratio $q\gtrsim 4$, even for modest initial eccentricities. We will also lose more than 10% of the signal information for the entire population of binary black holes that retain eccentricities $\gtrsim 0.2$ when they enter LIGO-Virgo-KAGRA band at 10Hz. These losses could amount to a $27\%$ loss in detection rate of high mass-ratio and/or moderate-to-high eccentric sources.
@@
Next we look at the change in optimal SNR [as defined in Eq. (14)] for binaries with eccentricities compared to those without. This will help understand how much further will our current and future GW detectors be able to see eccentric binary sources compared to quasi-circular ones. In Fig. 10 we show the relative change in optimal SNR for binaries as a function of initial eccentricity $e_{10}$. Each panel also shows contours (in red) of the actual optimal SNR values. In the top row, the three panels correspond to aLIGO, ET and CE sensitivities and to $20M_{\odot}$ binaries. The bottom row is similarly arranged, but corresponds to more massive $60M_{\odot}$ sources. The top left panel shows that eccentricity has a minimal effect in the overall loudness of signals coming from $20M_{\odot}$ BBHs as apparent to the current generation detectors. The panel below it shows that the same is true for more massive binaries as well. The top middle panel shows that for the third generation detector Einstein Telescope we can gain up to $3\%$ in overall signal strength compared to quasi-circular sources, corresponding to about $10\%$ larger detection volume. In the lower middle panel, we see a similar trend for more massive binaries as well, with the most eccentric sources recording $6\%$ larger SNRs for the most eccentric sources. Note that these binaries have initial eccentricities defined at $5$Hz instead of $10$Hz, as will be for the next generation detectors, because they are projected to be sensitive to lower frequencies than current-generation instruments. Because of this, our sources in the middle and right columns will have lower $e_{10}$ by the time they enter the LIGO sensitive band than those in the left column. Finally in the right top and bottom panels, we see that for Cosmic Explorer as well, eccentric sources will be up to $2-4\%$ louder than the quasi-circular ones. From these results it is clear that the horizon range will be $10-20\%$ larger for eccentric sources than quasi-circular ones for next generation detectors, increasing the likelihood of our finding a large population of dynamical capture binaries.
@@
To assess the effect of the subdominant modes compared to the dominant (quadrupole) mode for eccentric binary sources, in Fig. 11, we show the difference in optimal SNR with and without the inclusion of subdominant modes in eccentric waveform templates. We show this as a function of eccentricity $e$ and mass ratio $q$ in each panel, for LIGO (in the left column), ET (middle column), and CE (right column) detectors with total mass fixed to $20M_{\odot}$ (in the top row) and $60M_{\odot}$ (bottom row), dimensionless spins $(0.6,0.6)$ and source inclination angles of $\pi/3$ radians. As above, we keep the lower frequency cutoff $f_{\rm{low}}=10\rm{Hz}$ for LIGO and $5$Hz for CE and ET. We immediately note that the overall signal strength (optimal SNR) is higher by up to $3\%$ for high total mass binaries of $60M_{\odot}$ when we include sub-dominant modes. The same remains below $1\%$ for $20M_{\odot}$ binaries. This indicates that similar to quasi-circular binaries, higher-order modes are strongly excited close to merger than during earlier inspiral cycles for eccentric binaries as well. While this does not seem to motivate the use of higher-order modes in templates very strongly, it mostly implies that the overall signal amplitude is minimally affected by them. In the next study where we study the detectability of eccentric sources, which is a lot more sensitive to signal phasing than amplitude, we will find that indeed sub-dominant modes can lead to a non-trivial loss in recovered SNR for high mass-ratio systems.
@@ VI Conclusions and Future Outlook
Orbital eccentricity is a unique signature of compact binaries formed in dense stellar environments via dynamical processes. Such binaries are expected to comprise an important sub-population of all signals that ground-based GW detectors are expected to observe. Contemporary GW searches (and signal analyses in general) predominantly rely on waveform models tailored to quasi-circular binaries, primarily because quasi-circular binaries are expected to remain a majority of all detected GW signals, and also because the dimensionality of their parameter space is more manageable. Theoretical modeling of eccentric sources has also been historically ignored in favor of modeling non-eccentric sources including all effects of component spin angular momenta. However with more than a hundred signals having been already seen during the first three observation runs of the LIGO-Virgo-KAGRA instruments, we have clearly entered a signal dominated era. It is now timely to target astrophysically interesting sub-populations of compact binaries, such as eccentric BBHs. Many waveform models have also been under active development in the past few years catering to eccentric orbits [116, 117, 118, 100, 118, 119, 120, 121, 104, 122, 123, 124, 1, 60, 125, 92]. In this paper, we introduce ESIGMAHM, a time domain IMR waveform model for binaries of spinning black-holes on eccentric orbits, including the effect of higher-order modes. This model builds on the ENIGMA framework that was the very first IMR waveform model for eccentric binaries [122].
@@
ESIGMAHM is built of two pieces - an inspiral piece attached to a MR model. The inspiral segment is built upon the results coming from the post-Newtonian (PN) [146], self-force approach and black hole perturbation theory (BHPT) developed in the last several decades. The gravitational wave strain is obtained from spherical harmonic modes following Eq. (7). The instantaneous contribution to modes including the effect of spin and eccentricity are included up to 3.5PN order [97, 106, 107]. Additionally, the spinning and non-spinning hereditary contributions for quasi-circular orbits are included up to 3.5PN order following [158]. Recently reported [160, 169] quasi-circular non-spinning 4PN piece in the $\ell=|m|=2$ modes is also included in our framework. The conservative and radiative dynamics of the system is governed by the evolution equations $\dot{\phi}$, $\dot{l}$, and $\dot{x}$, $\dot{e}$, respectively. Though the effect of spins and eccentricity are included up to 3PN in evolution equations, the effect of spins in quasi-circular orbits is updated up to 4PN in $\dot{x}$. Besides, the 4PN non-spinning quasi-circular piece is also included in $\dot{x}$ following [169, 160]. The 3.5PN cubic-in-spin, 4PN non-spinning quasi-circular pieces in $\dot{x}$ are new and we list them in Appendix B. For the MR part, we inherit the approximation of ENIGMA that the binary has circularized (i.e., it has radiated away its eccentricity via GW emission) by the time it has entered the sensitivity band of the ground-based detectors. In Fig. 1 the time evolution of orbital angular velocity and the amplitude of the eccentric spinning (grey lines) and quasi-circular non-spinning (brown line), NR simulations are shown. It can be observed that at the transition from late inspiral to MR (denoted by filled circles) most systems have decayed nearly all their eccentricity. Following this approximation, we use NRSur7dq4 [91] - a state-of-the-art spinning quasi-circular surrogate model - to provide the MR piece. It is important to note that our approximation is shared by most contemporary modeling efforts for eccentric binaries.
@@
Having presented an upgraded IMR, spinning, eccentric waveform model with higher-order gravitational wave modes here, we plan to improve the same by comparison and calibration to a larger population of eccentric NR simulations in future work. Additionally, ESIGMA currently allows for component spins that are parallel or anti-parallel to orbital angular momentum. We plan to include the effect of arbitrary orientation of spins about orbital angular momentum (spin precession) in future work.
@@ VII Acknowledgements
We thank K. G. Arun and Guillaume Faye for discussions and insights into the modeling of spinning compact binaries. We are also thankful to the SXS Collaboration for making public a catalog of numerical relativity waveforms. K.P. thanks members of the Gravitation and Cosmology group at the Department of Physics, Indian Institute of Technology (IIT) Madras, for useful discussions. C.K.M acknowledges the hospitality of Institut d’Astrophysique de Paris and Max Planck Institute for Gravitational Physics (Albert Einstein Institute, Potsdam) during the final stages of writing of the paper. K.P. thanks the International Centre for Theoretical Sciences (ICTS) for the hospitality during the academic visits. A.M. and P.K. acknowledge support of the Department of Atomic Energy, Government of India, under project no. RTI4001 at the ICTS. P.K. also acknowledges support by the Ashok and Gita Vaish Early Career Faculty Fellowship at ICTS. C.K.M. acknowledges the support of SERB’s Core Research Grant No. CRG/2022/007959. Calculations in this paper were performed on the Sonic computing cluster at ICTS, and the Powehi workstation at IIT Madras. The authors are also grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. This document has LIGO preprint number LIGO-P2400383. We thank Antoni Ramos-Buades for useful comments and suggestions on our manuscript.

### v2

@@ V Importance of including eccentricity and higher-order waveform modes
An important effect that immediately distinguishes an eccentric binary from a quasi-circular one is the increased rate of inspiral, leading to a shorter time to merger. We can quantify this by computing the number of cycles $\rm{N}_{\rm{GW}}$ present in the dominant GW mode from a given (orbit-averaged) frequency to merger, for both eccentric ($\rm{N}^{\rm{ecc}}_{\rm{GW}}$) and equivalent quasi-circular ($\rm{N}^{\rm{circ}}_{\rm{GW}}$) cases with the same black hole masses and spins. We compute the differences between them
@@
starting from a (orbit-averaged) GW frequency of $10$Hz. Positive values of $\Delta\rm{N}_{\rm{GW}}$ imply a shorter eccentric signal. We show $\Delta\rm{N}_{\rm{GW}}$ in Fig. 9 as a function of total mass and mass ratio $M-q$ (left panel), and eccentricity and mass ratio $e-q$ (middle and right panels). The component spins are fixed to moderate values of $(0.5,0.5)$ in all panels, and orbital eccentricity is held constant at $0.05$ in the left panel. The middle panels correspond to a fixed binary mass of $20M_{\odot}$, while the right panel is for heavier binaries with mass $60M_{\odot}$. In the left panel showing $M-q$ plane, we see that there is a sharp increase in $\Delta\rm{N}_{\rm{GW}}$ with increasing mass ratio for small total masses, reaching as high as $(\sim 37.1)$ fewer GW cycles for $10M_{\odot}$ eccentric binaries with mass ratios $q\simeq 8$ even for relatively small value of eccentricity at $10$Hz $(e_{10})$ of $0.05$. For masses larger than $20M_{\odot}$ the difference is relatively smaller $(\lesssim 4)$ irrespective of mass ratio. The mass dependence is also relatively weak indicating that binaries starting with $e_{10}=0.05$ with masses above $20M_{\odot}$ rapidly circularize in the band of ground-based detectors. In $e-q$ plane shown in the middle panel, we see clearly a stronger non-linear growth in $\Delta\rm{N}_{\rm{GW}}$ as we increase initial eccentricity or (to a lesser extent) the binary mass ratio. For these $20M_{\odot}$ binaries, $\Delta\rm{N}_{\rm{GW}}$ grows rapidly and reaches $\mathcal{O}(10^{2})$ cycles when eccentricities is larger than $0.25$ for $q=2$ binaries, and when $e_{10}>0.15$ for $q=6$ systems. Compare this panel with the right panel to note the effect of increasing binary mass from $20$ to $60M_{\odot}$, a range which spans the masses of a bulk of sources observed by the current generation detectors so far [6, 7, 9, 220]. At $60M_{\odot}$ we notice a smaller yet substantial change in the number of GW cycles from eccentric binaries compared to quasi-circular ones. This is consistent with there being much fewer signals cycles in frequency band for higher masses. Note that the left and middle panels explore different parts of the parameter space, with only a horizontal strip of constant $M=20M_{\odot}$ in the left panel being common with the $e_{10}=0.05$ horizontal strip in the middle panel.
@@
Figure 12: Effect of eccentricity and higher-order modes are shown here on matched-filtering based GW searches. We show the distribution of fitting factors for different kinds of signal populations (as labelled in the legend), against a template bank of quasi-circular aligned-spin BBH waveform templates, similar to what is used in contemporary GW searches in LIGO-Virgo-KAGRA data. Comparing the blue and green curves shows the effectiveness of the same template bank in recovering eccentric vs quasi-circular BBH merger GW signals. This shows how currently used template banks will lose a sizable fraction of signal information from eccentric sources. The effect gets worse when sub-dominant harmonics are included in signal waveforms (comparing light and dark blue curves).
@@
Figure 13: Effect of eccentricity and higher-order waveform modes are shown on the detectability of eccentric binary mergers by current LIGO-Virgo-KAGRA searches. We show fitting factors for eccentric dominant-mode signals in the top row, and for eccentric multi-modal signals in the bottom row. In each panel, the same is shown as a function of source eccentricity at $10$Hz, and either binary mass ratio (left column) or total mass (right column). Component masses are sampled uniformly between $5-50M_{\odot}$ and spins uniformly between $[-0.9,0.9]$. Even though fitting factor values go down to 0.7 or so (as seen in Fig. 12), we set the lower limit of these panels’ colorbars to 0.9 in order to improve the visual contrast and highlight those parameter space regions where fitting factors remained above 0.97. We immediately note that for realistic signals, i.e. the bottom row, our current GW searches that employ only dominant-mode aligned-spin quasi-circular waveform templates, we will lose more than $10\%$ of signal information (optimal SNR) for sources with mass-ratio $q\gtrsim 4$, even for modest initial eccentricities. We will also lose more than 10% of the signal information for the entire population of binary black holes that retain eccentricities $\gtrsim 0.2$ when they enter LIGO-Virgo-KAGRA band at 10Hz. These losses could amount to a $27\%$ loss in detection rate of high mass-ratio and/or moderate-to-high eccentric sources.
@@
Next we look at the change in optimal SNR [as defined in Eq. (14)] for binaries with eccentricities compared to those without. This will help understand how much further will our current and future GW detectors be able to see eccentric binary sources compared to quasi-circular ones. In Fig. 10 we show the relative change in optimal SNR for binaries as a function of initial eccentricity $e_{10}$. The relative change in SNR is defined as $\frac{\Delta\rho}{\rho_{0}}$, where $\Delta\rho=\rho_{0}-\rho_{\rm{ecc}}$ and $\rho_{0}$, $\rho_{\rm{ecc}}$ are the optimal SNR for circular and eccentric signals, respectively. Each panel also shows contours (in red) of the actual optimal SNR values. In the top row, the three panels correspond to aLIGO, ET and CE sensitivities and to $20M_{\odot}$ binaries. The bottom row is similarly arranged, but corresponds to more massive $60M_{\odot}$ sources. The top left panel shows that eccentricity has a minimal effect in the overall loudness of signals coming from $20M_{\odot}$ BBHs as apparent to the current generation detectors with a gain and loss in overall signal strength of up to $2.1\%$ and $0.6\%$ compared to quasi-circular sources, respectively. The panel below it shows that the same is true for more massive binaries as well with gain and loss of up to $8.2\%$ and $6.2\%$, respectively. The top middle panel shows that for the third generation detector Einstein Telescope we can gain up to $0.4\%$ and the loss can be $3.2\%$ in overall signal strength compared to quasi-circular sources of total mass $20M_{\odot}$. In the lower middle panel, we see a similar trend for more massive binaries as well, with the most eccentric sources recording $3\%$ larger SNRs for the most eccentric sources, corresponding to about $10\%$ larger detection volume. Note that these binaries have initial eccentricities defined at $5$Hz instead of $10$Hz, as will be for the next generation detectors, because they are projected to be sensitive to lower frequencies than current-generation instruments. Because of this, our sources in the middle and right columns will have lower $e_{10}$ by the time they enter the LIGO sensitive band than those in the left column. Finally in the right top and bottom panels, we see that for Cosmic Explorer as well, eccentric sources will be up to $2-4\%$ louder than the quasi-circular ones. Although we mostly gain SNR, we can also lose SNR depending on the mass value and the sensitivity of the detector. From these results it is clear that the horizon range will be $10-20\%$ larger for eccentric sources than quasi-circular ones for next generation detectors, increasing the likelihood of our finding a large population of dynamical capture binaries.
@@
To assess the effect of the subdominant modes compared to the dominant (quadrupole) mode for eccentric binary sources, in Fig. 11, we show the difference in optimal SNR with and without the inclusion of subdominant modes in eccentric waveform templates. We show this as a function of eccentricity $e$ and mass ratio $q$ in each panel, for LIGO (in the left column), ET (middle column), and CE (right column) detectors with total mass fixed to $20M_{\odot}$ (in the top row) and $60M_{\odot}$ (bottom row), dimensionless spins $(0.6,0.6)$ and source inclination angles of $\pi/3$ radians. As above, we keep the lower frequency cutoff $f_{\rm{low}}=10\rm{Hz}$ for LIGO and $5$Hz for CE and ET. We immediately note that the overall signal strength (optimal SNR) is higher by up to $3\%$ for high total mass binaries of $60M_{\odot}$ when we include sub-dominant modes. The same remains below $1\%$ for $20M_{\odot}$ binaries. This indicates that similar to quasi-circular binaries, higher-order modes are strongly excited close to merger than during earlier inspiral cycles for eccentric binaries as well. While this does not seem to motivate the use of higher-order modes in templates very strongly, it mostly implies that the overall signal amplitude is minimally affected by them. Next we study the detectability of eccentric sources, which is a lot more sensitive to signal phasing than amplitude, we will find that indeed sub-dominant modes can lead to a non-trivial loss in recovered SNR for high mass-ratio systems.
@@ VI Conclusions and Future Outlook
Orbital eccentricity is a unique signature of compact binaries formed in dense stellar environments via dynamical processes. Such binaries are expected to comprise an important sub-population of all signals that ground-based GW detectors are expected to observe. Contemporary GW searches (and signal analyses in general) predominantly rely on waveform models tailored to quasi-circular binaries, primarily because quasi-circular binaries are expected to remain a majority of all detected GW signals, and also because the dimensionality of their parameter space is more manageable. Theoretical IMR modeling of eccentric sources was largely ignored in literature in the pre-GW150914 detection era [2], in favor of modeling non-eccentric sources with very high precision.
However, since GW150914, more than 90 signals having been already seen during the first three observation runs of the LIGO-Virgo-KAGRA instruments, we have clearly entered a signal dominated era. It is now timely to target astrophysically interesting sub-populations of compact binaries, such as eccentric BBHs. Many waveform models have also been under active development in the past few years catering to eccentric orbits  [124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 130, 146, 147, 148, 134, 149, 150, 1, 151, 1, 152, 65, 62, 153, 101, 104, 154, 155, 156, 157, 158, 159, 93, 160, 161]. In this paper, we introduce ESIGMAHM, a time domain IMR waveform model for binaries of spinning black-holes on eccentric orbits, including the effect of higher-order modes. This model builds on the ENIGMA framework that was the very first IMR waveform model for eccentric binaries [149].
@@
ESIGMAHM is composed of two segments - an inspiral part attached to a PMR model. The inspiral segment is built upon the results coming from the post-Newtonian (PN) [179], self-force approach and black hole perturbation theory (BHPT) developed in the last several decades. The gravitational wave strain is obtained from spherical harmonic modes following Eq. (7). The instantaneous contribution to modes including the effect of spin and eccentricity are included up to 3.5PN order [127, 136, 138]. Additionally, the spinning and non-spinning hereditary contributions for quasi-circular orbits are included up to 3.5PN order following [191]. Recently reported [193, 203] quasi-circular non-spinning 4PN piece in the $\ell=|m|=2$ modes is also included in our framework. The conservative and radiative dynamics of the system is governed by the evolution equations $\dot{\phi}$, $\dot{l}$, and $\dot{x}$, $\dot{e}$, respectively. Though the effect of spins and eccentricity are included up to 3PN in evolution equations, the effect of spins in quasi-circular orbits is updated up to 4PN in $\dot{x}$. Besides, the 4PN non-spinning quasi-circular piece is also included in $\dot{x}$ following [203, 193]. The 3.5PN cubic-in-spin, 4PN non-spinning quasi-circular pieces in $\dot{x}$ are new and we list them in Appendix B. For the PMR part, we inherit the approximation of ENIGMA that the binary has circularized (i.e., it has radiated away its eccentricity via GW emission) by the time it has entered the sensitivity band of the ground-based detectors. In Fig. 1 the time evolution of orbital angular velocity and the amplitude of the eccentric spinning (grey lines) and quasi-circular non-spinning (brown line), NR simulations are shown. It can be observed that at the transition from late inspiral to PMR (denoted by filled circles) most systems have decayed nearly all their eccentricity. Following this approximation, we use NRSur7dq4 [92] - a state-of-the-art spinning quasi-circular surrogate model - to provide the PMR piece. [^9] EOB models of the SEOB and TEOB families attach a quasi-circular ringdown model at the peak of respective $(\ell,m)$ mode amplitudes. Their inspiral-merger piece is computed based on an eccentric dynamics. Smooth matching is achieved by applying next-to-quasi-circular corrections during the plunge. In ESIGMAHM, however, quasi-circular motion is assumed for a relatively larger portion of the signal, approximately starting from the Kerr or Schwarzschild ISCO.
@@
Having presented an upgraded IMR, spinning, eccentric waveform model with higher-order gravitational wave modes here, we plan to improve the same by comparison and calibration to a larger population of eccentric NR simulations in future work. Additionally, ESIGMA currently allows for component spins that are parallel or anti-parallel to orbital angular momentum. We plan to include the effect of arbitrary orientation of spins about orbital angular momentum (spin precession) in future work. Besides, the ESIGMA framework aims to iteratively improve waveform models for the eccentric inspiral and plunge stages, with the ultimate goal of utilizing eccentric NR surrogates to accurately describe the plunge-merger-ringdown.
@@ VII Acknowledgements
We thank K. G. Arun and Guillaume Faye for discussions and insights into the modelling of spinning compact binaries. We are also thankful to the SXS Collaboration for making public a catalog of numerical relativity waveforms. K.P. thanks members of the Gravitation and Cosmology group at the Department of Physics, Indian Institute of Technology (IIT) Madras, for useful discussions. C.K.M acknowledges the hospitality of Institut d’Astrophysique de Paris and Max Planck Institute for Gravitational Physics (Albert Einstein Institute, Potsdam) during the final stages of writing of the paper. K.P. thanks the International Centre for Theoretical Sciences (ICTS) for the hospitality during the academic visits. A.M. and P.K. acknowledge support of the Department of Atomic Energy, Government of India, under project no. RTI4001 at the ICTS. P.K. also acknowledges support by the Ashok and Gita Vaish Early Career Faculty Fellowship at ICTS. C.K.M. acknowledges the support of SERB’s Core Research Grant No. CRG/2022/007959. Calculations in this paper were performed on the Sonic computing cluster at ICTS, and the Powehi workstation at IIT Madras. The authors are also grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. This document has LIGO preprint number LIGO-P2400383. We thank Antoni Ramos-Buades for useful comments and suggestions on our manuscript. K. S. acknowledges support from the Long Term Visiting Students Program at ICTS during the course of this work.